hi guys i really need help on this problem:
Henry can ride his horse at 4 miles per hour and get to the battlefield on time. If he stops for 1 hour to m ake a speech to his troops, he must ride at 5 miles per hour for the whole trip to get to the battlefield on time. How far is it to the battlefield?
Can you please tell me step-by-step how to solve this because i have more of these kinds of problems on homework!
Thanks!

Answer by stanbon(75887)

Henry can ride his horse at 4 miles per hour and get to the battlefield on time. If he stops for 1 hour to m ake a speech to his troops, he must ride at 5 miles per hour for the whole trip to get to the battlefield on time. How far is it to the battlefield?
--------------
1st ride to the battlefield DATA:
Rate = 4 mph ; time = x hrs. ; distance = rt=4x miles
--------------
2nd ride to the battlefield DATA:
Rate = 5 mph ; time = x-1 hrs. ; distance = rt = 5(x-1) miles
-------------------
EQUATION:
1st distance = 2nd distance
4x=5(x-1)
4x=5x-5
-x=-5
x=5 hours
---------------
Distance = 4*5= 20 miles
Cheers,
Stan H.